# Multiscale Causal Learning

### v2

@@ 2.2.1 Learning and inference definitions:
(intuitive summary) To reiterate and summarise the above:
@@
Definition 5. [learning] .
  • A proxy $<$ is a binary relation on statements, and the set of all proxies is $Q$.
  • $<_{w}$ is the weakness proxy. For statements $l_{1},l_{2}$ we have $l_{1}<_{w}l_{2}$ iff $\lvert E_{l_{1}}\rvert<\lvert E_{l_{2}}\rvert$.
By the weakness of a statement, we mean the cardinality of its extension.
(generalisation) A statement $l$ generalises to a $\mathfrak{v}$-task $\alpha$ iff $l\in\Pi_{\alpha}$ ($l$ is a cause of $\alpha$). We speak of learning $\omega$ from $\alpha$ iff, given a proxy $<$, $\pi\in{\Pi}_{\alpha}$ maximises $<$ relative to all other policies in ${\Pi}_{\alpha}$, and $\pi\in{\Pi}_{\omega}$.
@@
(sample efficiency) Suppose $\mathfrak{app}$ is the set of all pairs of policies. Assume a proxy $<$ returns $1$ iff true, else 00. Proxy $<_{a}$ is more sample efficient than $<_{b}$ iff
@@
(optimal proxy) There is no proxy more sample efficient than $<_{w}$, so we call $<_{w}$ optimal. This formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [10]).
@@
(intuitive summary) Learning is an activity undertaken by some manner of intelligent agent, and a task has been “learned” by an agent that knows a correct policy. Humans typically learn from “examples”. An example of a task is a correct output and input. A collection of examples is a child task, so “learning” is an attempt to generalise from a child, to one of its parents. The lower level the child from which an agent generalises to parent, the “faster” it learns, the more sample efficient the proxy. The most sample efficient proxy is $<_{w}$.
@@
Definition 6. [uninstantiated-tasks] The set of all tasks with no abstraction (meaning $\mathfrak{v}=P$) is $\Gamma_{P}$ (it contains every task in every vocabulary). For every $P$-task $\rho\in\Gamma_{P}$ there exists a function $\lambda_{\rho}:2^{P}\rightarrow\Gamma_{P}$ that takes a vocabulary $\mathfrak{v}^{\prime}\in 2^{P}$ and returns a highest level child $\omega\sqsubset\rho$ which is also a $\mathfrak{v}^{\prime}$-task. We call $\lambda_{\rho}$ an uninstantiated-task, and $\lambda_{1}\sqsubset\lambda_{2}$ iff $\lambda_{1}(P)\sqsubset\lambda_{2}(P)$.
@@
Definition 7. [utility of intelligence] Utility is the difference in weakness between the weakest and strongest correct policies of a task. The utility of a $\mathfrak{v}$-task $\gamma$ is $\epsilon(\gamma)=\underset{{\pi}\in{\Pi}_{\gamma}}{\max}\left(\lvert E_{\pi}\rvert-\lvert{O}_{\gamma}\rvert\right)$.
@@
(intuitive summary) Uninstantiated tasks let us represent the same task in different abstraction layers. We can then use utility to measure how “good” an abstraction layer is for a particular task. This recognises that abstraction layers are not value-neutral. Each is better or worse for some tasks. Only the absence of abstraction is entirely neutral.
3 Arguments and Results
We must first explain how $\mathfrak{v}$-tasks can represent MCAs. We must explain how different layers of the MCA are formed; how “the behaviour of subunits percolates up” to determine higher level behaviour, which in turn “constrains and facilitates” lower level behaviours [26]. Once that is done we show how an MCA can facilitate more efficient WPO than a static ISA. We conceive of an MCA as a causal ‘chain’ where each node is the goal directed behaviour of the node that preceded it. For example, cells form organs which form organisms.
@@ 3 Arguments and Results
We make no distinction between an object and its behaviour. Embodied and embedded, behaviour is limited to a particular vocabulary $\mathfrak{v}$ (meaning there the number of possible configurations is finite). The behaviour of two cells as individuals is two $\mathfrak{v}$-tasks $\alpha_{1}$ and $\alpha_{2}$. The collective behaviour of the two cells is a third $\mathfrak{v}$-task $\omega$ s.t. $\alpha_{1}\sqsubset\omega$ and $\alpha_{2}\sqsubset\omega$. Intuitively, this just means the two cells are parts of a larger whole. If there exists a correct policy for $\omega$, then the cells can behave in service of a common goal and thus have a common identity (for example if both cells use policies in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$). Note this identity is a precursor to the “causal identity for self” [11] that differentiates observation from intervention in the manner of reafference [33].
@@
If the parts of a larger whole learn, then the probability that they learn a shared policy is maximised if they use the weakness proxy (what we have been referring to as WPO). This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks [10, prop. 1,2]. For our cell example, a sufficiently weak policy at the level of individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$. In other words WPO is not just facilitated by an MCA, but required for high level behaviour. To explain how higher levels constrain lower, consider that higher levels in an MCA are just an abstraction; something we ascribe to the behaviour of the lowest level. For the sake of intuition; different levels are patterns within one larger pattern, and we simplify the larger pattern by dividing it up. For example, a collective of cells remains cells whether we call it an organ or not. A collective does not exist in the absence of its parts. $\omega$ and $\alpha_{1}$ and $\alpha_{2}$ all exist in the same vocabulary. If the outputs of $\alpha_{1}$ affect the inputs of $\alpha_{2}$ and vice versa, then the outputs of $\omega$ determine the inputs of $\alpha_{1}$ and $\alpha_{2}$. The collective constrains behaviour at the lower level. We point this out to show how a $\mathfrak{v}$-task can represent an MCA. Just as the weakness proxy [10, proposition 1,2] facilitates causal learning at a single scale [11], in the context of an MCA it will factilitate causal learning at multiple scales, because the MCA as a whole is just another $\mathfrak{v}$-task.

### v3

@@ 2.1 Learning and inference definitions
To reiterate and summarise the above:
@@
{adjustwidth}
11.6pt11.6pt
Definition 5 (learning).
A proxy $<$ is a binary relation on statements. In this paper we use only one proxy, called the weakness proxy, which compares the cardinality of a statement’s extension. For statements $l_{1},l_{2}$ we have $l_{1}<l_{2}$ iff $\lvert Z_{l_{1}}\rvert<\lvert Z_{l_{2}}\rvert$. Whenever we use $<$ to compare statements, we are referring to the aforementioned weakness proxy.
(generalisation) A statement $l$ generalises to a $\mathfrak{v}$-task $\alpha$ iff $l\in\Pi_{\alpha}$. We speak of learning $\omega$ from $\alpha$ iff, given a proxy $<$, $\pi\in{\Pi}_{\alpha}$ maximises $<$ relative to all other policies in ${\Pi}_{\alpha}$, and $\pi\in{\Pi}_{\omega}$.
@@
(efficiency) Suppose $\mathfrak{app}$ is the set of all pairs of policies. Assume a proxy $<$ returns $1$ iff true, else 00. Proxy $<_{a}$ is more efficient than $<_{b}$ iff
@@
(optimal proxy) There is no proxy more efficient than weakness. The weakness proxy formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [34]).
@@
Learning is an activity undertaken by some manner of intelligent agent, and a task has been “learned” by an agent that knows a correct policy. Humans typically learn from “examples”. An example of a task is a correct output and input. A collection of examples is a child task, so “learning” is an attempt to generalise from a child to one of its parents. The lower level the child from which an agent generalises to parent, the “faster” it learns (it chooses policies that complete a wider variety of tasks, and thus are more sample and energy efficient choices), the more efficient the proxy. The most efficient proxy is weakness (see proofs 1 and 2, or [34]), which is why we’re using it here.
{adjustwidth}
11.6pt11.6pt
Definition 6 (uninstantiated-tasks).
@@
The set of all tasks with no abstraction (meaning $\mathfrak{v}=P$) is $\Gamma_{P}$ (it contains every task in every vocabulary). For every $P$-task $\rho\in\Gamma_{P}$ there exists a function $\lambda_{\rho}:2^{P}\rightarrow\Gamma_{P}$ that takes a vocabulary $\mathfrak{v}^{\prime}\in 2^{P}$ and returns a highest level child $\omega\sqsubset\rho$ which is also a $\mathfrak{v}^{\prime}$-task. We call $\lambda_{\rho}$ an uninstantiated-task, and $\lambda_{1}\sqsubset\lambda_{2}$ iff $\lambda_{1}(P)\sqsubset\lambda_{2}(P)$.
Definition 7 (utility of intelligence).
@@
Utility is the difference in weakness between the weakest and strongest correct policies of a task. The utility of a $\mathfrak{v}$-task $\gamma$ is $\epsilon(\gamma)=\underset{{\pi}\in{\Pi}_{\gamma}}{\max}\left(\lvert E_{\pi}\rvert-\lvert{O}_{\gamma}\rvert\right)$.
@@
Uninstantiated tasks let us represent the same task in different abstraction layers. We can then use utility to measure how “good” an abstraction layer is for a particular task [12]. This recognises that abstraction layers are not value-neutral. Each is better or worse for some tasks. Only the absence of abstraction is entirely neutral.
2.2 Formalising Multilayer, Multiscale Architectures
The following definition is explained at length over the remainder of this paper. Refer back here as needed.
{adjustwidth}
11.6pt11.6pt
Definition 8 (multiscale competency architecture).
  • The causal chain of layers of increasing abstraction is a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$.
  • $\mathfrak{f}:2^{L_{\mathfrak{v}}}\rightarrow 2^{P}$ is a function that takes the extension $E_{\pi}$ of a policy $\pi\in L_{\mathfrak{v}}$, and returns a vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$.
  • The state of the MCA is a sequence of policies $\langle\pi^{0},\pi^{1}...\pi^{n}\rangle$ and a sequence of vocabularies $\langle\mathfrak{v}^{0},\mathfrak{v}^{1}...\mathfrak{v}^{n}\rangle$ such that $\mathfrak{v}^{i+1}=\mathfrak{f}(E_{\pi^{i}})$ and $\pi^{i}\in\Pi_{\lambda^{i}(\mathfrak{v^{i}})}$.
In the absence of abstraction where the system is seen as nothing more than the sum of its parts, the MCA is just a task $\lambda^{0}(\mathfrak{v}^{0})$, allowing us to look at the system across scales of distribution. We say the MCA is over-constrained when there exists $i<n$ s.t. and $\Pi_{\lambda^{i}(\mathfrak{v^{i}})}=\emptyset$, and multilayer-causal-learning (MCL) occurs when the MCA is under-constrained and the proxy for learning is weakness. Note that the vocabulary is different at each level in the MCA, which means each has its own generational hierarchy of tasks. By a higher level of abstraction in the MCA, we mean a later task in the causal chain.
2.3 Proofs
Here we present three proofs. The first two demonstrate the optimality of the weakness proxy, and the last show that adaptability later in the causal chain necessitates adaptability earlier in the chain. Preliminary proof and experimental results showing weakness to be the optimal proxy were published in conference proceedings [34]. However, the proofs contained minor errors. Most notably the earlier proof of sufficiency failed to account for task equivalency and so miscounted the total number of tasks, rather than counting the total number of equivalence classes [^11] Credit goes to Nora Belrose for spotting this error.. This does not change the result, but for the sake of rigour these errors must be addressed. To that end, here we present more detailed alternatives to those proofs, which once again show the optimality of weakness as a proxy.
Proposition 1 (sufficiency).
Assume $\alpha\sqsubset\omega$. The weakness proxy sufficient to maximise the probability that a parent $\omega$ is learned from a child $\alpha$ [^12] Assume there exist correct policies for $\omega$, because otherwise there would be no point in trying to learn it..
Proof.
You’re given the definition of $\mathfrak{v}$-task $\alpha$ from which you infer a hypothesis $\pi\in{\Pi}_{\alpha}$. To learn $\omega$, you need $\pi\in\Pi_{\omega}$:
  1. 1.
    For every $\pi\in\Pi_{\alpha}$ there exists a $\mathfrak{v}$-task $\gamma_{\pi}\in\Gamma_{\mathfrak{v}}$ s.t. $O_{\gamma_{\pi}}=E_{\pi}$, meaning $\pi$ permits only correct outputs for that task regardless of input. We’ll call the highest level task $\gamma_{\pi}$ s.t. $O_{\gamma_{\pi}}=E_{\pi}$ the policy task of $\pi$.
  2. 2.
    $\omega$ is either the policy task of a policy in $\Pi_{\alpha}$, or a child thereof.
  3. 3.
    If a policy $\pi$ is correct for a parent of $\omega$, then it is also correct for $\omega$. Hence we should choose $\pi$ that has a policy task with the largest number of children. As tasks are uniformly distributed, that will maximise the probability that $\omega$ is $\gamma_{\pi}$ or a child thereof.
  4. 4.
    For the purpose of this proof, we say one task is equivalent [^13] This is because switching from $\beta$ to $\zeta$ s.t. $I_{\beta}\neq I_{\zeta}$ and $O_{\beta}=O_{\zeta}$ would be to pursue the same goal in different circumstances. This is because inputs are subsets of outputs, so both sets of inputs are implied by the outputs. $O_{\zeta}$ implies $I_{\beta}$ and $O_{\beta}$ implies $I_{\zeta}$ to another if it has the same correct outputs.
  5. 5.
    No two policies in $\Pi_{\alpha}$ have the same policy task [^14] Every policy task for policies of $\alpha$ is non-equivalent from the others.. This is because all the policies in $\Pi_{\alpha}$ are derived from the same set inputs, $I_{\alpha}$.
  6. 6.
    The set of statements which might be outputs addressing inputs in ${I}_{\omega}$ and not ${I}_{\alpha}$, is $\overline{E_{{I}_{\alpha}}}=\{l\in L_{\mathfrak{v}}:l\notin E_{{I}_{\alpha}}\}$ [^15] This is because $E_{{I}_{\alpha}}$ contains every statement which is a correct output or an incorrect output, and $\overline{E_{{I}_{\alpha}}}$ contains every statement which could possibly be in $I_{\omega}$, $E_{I_{\omega}}$ and thus $O_{\omega}$..
  7. 7.
    For any given $\pi\in{\Pi}_{\alpha}$, the extension $E_{\pi}$ of $\pi$ is the set of outputs $\pi$ implies. The subset of $E_{\pi}$ which fall outside the scope of what is required for the known task $\alpha$ is $\overline{E_{{I}_{\alpha}}}\cap E_{\pi}$ [^16] This is because $E_{{I}_{\alpha}}$ is the set of all conceivable outputs by which one might attempt to complete $\alpha$, and so the set of all outputs that can’t be made when undertaking $\alpha$ is $\overline{E_{{I}_{\alpha}}}$ because those outputs occur given inputs that aren’t part of ${I}_{\alpha}$..
  8. 8.
    $\left|\overline{E_{{I}_{\alpha}}}\cap E_{\pi}\right|$ increases monotonically with $\left|E_{\pi}\right|$, because for all $e\in E_{\pi}$, $e\notin\overline{E_{{I}_{\alpha}}}$ iff $e\in E_{{I}_{\alpha}}$. This is because $L_{\mathfrak{v}}=\overline{E_{{I}_{\alpha}}}\cup E_{{I}_{\alpha}}$ and for all $\pi\in\Pi_{\alpha}$, $E_{\pi}\subseteq L_{\mathfrak{v}}$.
  9. 9.
    $2^{\left|\overline{E_{{I}_{\alpha}}}\cap E_{\pi}\right|}$ is the number of non-equivalent parents of $\alpha$ to which $\pi$ generalises. It increases monotonically with the weakness of $\pi$.
  10. 10.
    Given $\mathfrak{v}$-tasks are uniformly distributed and $\Pi_{\alpha}\cap\Pi_{\omega}\neq\emptyset$, the probability that $\pi\in{\Pi}_{\alpha}$ generalises to $\omega$ is
    $p(\pi\in{\Pi}_{\omega}\mid\pi\in{\Pi}_{\alpha},\alpha\sqsubset\omega)=\frac{2^{\left|\overline{E_{{I}_{\alpha}}}\cap E_{\pi}\right|}}{2^{\left|\overline{E_{{I}_{\alpha}}}\right|}}$
$p(\pi\in{\Pi}_{\omega}\mid\pi\in{\Pi}_{\alpha},\alpha\sqsubset\omega)$ is maximised when $\left|E_{\pi}\right|$ is maximised. Recall from definition 4 that $<_{w}$ is the weakness proxy. For statements $l_{1},l_{2}$ we have $l_{1}<_{w}l_{2}$ iff $\lvert E_{l_{1}}\rvert<\lvert E_{l_{2}}\rvert$. $\pi$ that maximises $<_{w}$ will also maximise $p(\pi\in{\Pi}_{\omega}\mid\pi\in{\Pi}_{\alpha},\alpha\sqsubset\omega)$. Hence the weakness proxy maximises the probability that [^17] Subsequently it also maximises the efficiency with which a parent $\omega$ is learned from a child $\alpha$. a parent $\omega$ is learned from a child $\alpha$. ∎
Proposition 2 (necessity).
To maximise the probability of learning $\omega$ from $\alpha$, it is necessary to use weakness as a proxy.
Proof.
Let $\alpha$ and $\omega$ be defined exactly as they were in the proof of prop. 1.
  1. 1.
    If $\pi\in{\Pi}_{\alpha}$ and $E_{{I}_{\omega}}\cap E_{\pi}={O}_{\omega}$, then it must be he case that ${O}_{\omega}\subseteq E_{\pi}$.
  2. 2.
    If $\left|E_{\pi}\right|<\left|{O}_{\omega}\right|$ then generalisation cannot occur, because that would mean that ${O}_{\omega}\not\subseteq E_{\pi}$.
  3. 3.
    Therefore generalisation is only possible if $\left|E_{\pi}\right|\geq\left|{O}_{\omega}\right|$, meaning a sufficiently weak hypothesis is necessary to generalise from child to parent.
  4. 4.
    For any two hypotheses $\pi_{1}$ and $\pi_{2}$, if $\mid E_{\pi_{1}}\mid<\mid E_{\pi_{2}}\mid$ then the probability $p(\left|E_{\pi_{1}}\right|\geq\left|{O}_{\omega}\right|)<p(\left|E_{\pi_{2}}\right|\geq\left|{O}_{\omega}\right|)$ because tasks are uniformly distributed.
  5. 5.
    Hence the probability that $\left|E_{{m}}\right|\geq\left|{O}_{\omega}\right|$ is maximised when $\left|E_{{m}}\right|$ is maximised. To maximise the probability of learning $\omega$ from $\alpha$, it is necessary to select the weakest hypothesis.
To select the weakest hypothesis, it is necessary to use the weakness proxy. ∎
Proposition 3.
The greater the utility $\epsilon(\lambda^{i+1}(\mathfrak{v}^{i+1}))$, the weaker the policy $\pi^{i}$ s.t. $\mathfrak{f}(E_{\pi^{i}})=\mathfrak{v}^{i+1}$ must be.
Proof.
3 Multiscale, Multilayer Causal Learning
To begin, we must formally describe the architecture before we can consider its strengths and weaknesses in comparison to a typical computer. Hence we describe how $\mathfrak{v}$-tasks can describe the multiple scales and layers of abstraction in a biological system. We frame this first by looking at distribution of control within the system, and second by formalising the levels of abstraction at which we can view the system as a whole. Intuitively, this means we formalise different cross sections of the system.
Looking at how control is distributed across different scales within the system we can see how the individual parts that make up a collective each possess their own policy or “identity”, and how that will (or will not) align with a shared collective policy or identity. This treats the system as nothing more than the sum of its parts. It is about self organisation, and how control is distributed within the resulting system. For example, individual neurons in the brain engage in predictive processing of eachother, and emulating this approach has yielded promising results in artificial neural networks [25, 26, 27]. They form a network with a single coherent policy, and yet at the same time are individual neurons engaging in active inference.
In contrast abstraction is about how the behaviour of cells enact an organ, and the behaviour of organs enacts an organism which enacts even more abstract behaviours like, for example, language. Each of these is a layer of abstraction. Layers exist in a causal ‘chain’ (a directed acyclic graph) where each layer (node) is the goal directed behaviour of the layer that preceded it. For example, cells form organs which form organisms [45].
@@ 3 Multiscale, Multilayer Causal Learning
3.1 Distributing Self Organisation Across Scales
The formalism we extend is enactive [12, 30], describing a self-organising system by describing what it does, rather than what it is. An embodied, embedded system can only do so many things, and so its behaviour is expressed using a finite vocabulary $\mathfrak{v}$ of declarative programs. These programs describe everything the system can do. For example if a system can be in a state of pain, then there are programs in $\mathfrak{v}$ which are “true” when the system is in such a state. Since states of the environment are mutually exclusive, $\mathfrak{v}$ implies an embodied formal language. By language here, we do not mean like the words humans use to communicate. We mean a language of concrete computations [52], where every statement in that language is rather than describes part of the state of a system (put another way, the interpreter is not assumed to be a UTM, but physical laws). For example, consider the language implied by a human being. If the human raises their arm, then they are expressing a statement in their embodied language. That statement then constrains what the body can do next, because everything that happens next begins from the point where the person had their arm raised [30].
We frame self organisation in the same terms as the computer analogy we employed in the introduction. The behaviour of a self-organising system is goal directed, meaning there exists a constraint on its behaviour that defines that goal. The behaviour that satisfies the goal is expressed in the embodied formal language; it is a set of inputs and outputs we call a $\mathfrak{v}$-task. The inputs and outputs in the $\mathfrak{v}$-task are ‘fit’ or ‘correct’, the rest are not. Given these inputs, the system is fit if it embodies a statement that constrains behaviour to correct outputs only. We call statements that constrain behaviour policies. Policies constrain outputs given inputs, as in the computer analogy. Just as many different neural topologies can have the same functionality, many different policies can imply the same set of outputs given a particular set of inputs. A system learns if it can construct a policy from historical data which constrains behaviour to correct outputs (generalises beyond historical data), and once again the optimal choice of policy is the weakest. For brevity, when a system optimises for weak policies, we say it is engaged in weak policy optimisation (WPO).
The above vocabulary can describe a distributed system. For example, an organ $\omega$ formed of two cells $\alpha_{1}$ and $\alpha_{2}$. The behaviour of two cells as individuals is two $\mathfrak{v}$-tasks $\alpha_{1}$ and $\alpha_{2}$. The collective behaviour of the two cells is a third $\mathfrak{v}$-task $\omega$ s.t. $\alpha_{1}\sqsubset\omega$ and $\alpha_{2}\sqsubset\omega$. Intuitively, this just means the two cells are parts of a larger whole. If there exists a correct policy for $\omega$, then the cells can behave in service of a common goal and thus have a common identity (for example if both cells use policies in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$). Note this identity is a precursor to the “causal identity for self” [16] that, in its most basic form, differentiates observation from intervention in the manner of reafference [11, 30].
@@ 3.1 Distributing Self Organisation Across Scales
A policy is like an identity described in the MCA literature [45]. The probability that the parts of a system embody a collective policy is higher if they embody weaker policies. This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks (see proofs 1, 2 or [34]).
Consider the aforementioned $\alpha_{1}$ and $\alpha_{2}$. A sufficiently weak policy at the level of individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$.
A collective of cells remains cells regardless of whether we deem it to be an organ. The parts are still there. Formally the cells $\alpha_{1}$ and $\alpha_{2}$ and the collective $\omega$ share a vocabulary. Just as the outputs of $\alpha_{1}$ can affect the inputs of $\alpha_{2}$ and vice versa, the outputs of $\omega$ determine the inputs of $\alpha_{1}$ and $\alpha_{2}$. The difference is a matter of scale. Just as the weakness proxy [34, proposition 1,2] facilitates causal learning at a single scale [16], in the context of an MCA it will factilitate causal learning at multiple scales, because the MCA as a whole is just another $\mathfrak{v}$-task. However, this still just treats everything as a collective of cells.
3.2 The Causal Chain of Abstraction
From the enactive perspective, a human is something cells do, rather than cells are.
To further clarify, we call a system complex when its behaviour is something we cannot easily comprehend or predict. However, just because we cannot comprehend and predict all of a system’s behaviour does not mean we cannot predict some of it. We suggest that so-called emergent properties are merely those aspects of a complex system’s behaviour which we can comprehend and predict. Aspects and patterns within patterns that remain comprehensible. For example, the overall pattern of a fractal like the Barnsley Fern [53] can be obvious, while the details elude comprehension.
We now illustrate how behaviour forms a causal chain of layers of increasingly abstract behaviour, from cells to organs to organisms to the behaviour of those organisms (specifically human abstractions). We’ll then show how it is possible to ’intervene’ in that causal chain to disconnect more abstract behaviour from the less abstract behaviour that gave rise to it. We show this reduces “utility” (def. 7) for tasks at higher levels, meaning reduced capacity for WPO. This is because the resulting distribution of control is altered when we intervene in the causal chain.
Just as an organ boils down to a collection of cells, the collective of cells as a whole is an abstraction layer in which “organ” runs, like software. However, software is just a state of hardware [12]. Put another way, “organ” is a classification we ascribe to the the behaviour of a cellular collective as a whole.
As stated in the introduction, we’re interested in how MCAs relate to abstraction layers in a computer. x86 is a classification we ascribe to the product of human behaviour, a CPU. It is a higher level behaviour of an MCA that includes humans. To understand how we might “intervene” in the causal chain and disconnect higher level behaviours from the system which gave rise to them, we need to abstract away the lower levels. To relate this back to the cell example, we must abstract away individual cells and focus on the consequences of their collective behaviour, as a unified whole.